# Advances in 4D Representation: Geometry, Motion, and Interaction

### v1

@@ 3.2.1 Rigid Motion
Rigid bodies include manufactured items (boxes, chairs, kitchenware) and vehicles, which are often approximated as rigid despite minor internal deformations [HLS24].
@@ 3.2.2 Non-Rigid Motion
where $\Delta_{\theta}$ predicts displacements conditioned on temporal code $\tau(t)$. These fields are trained from 2D observations via differentiable rendering, eliminating 3D supervision. However, sparse views complicates learning, requiring regularization for robustness.
@@ 3.2.3 Integrating physical priors
Physics-based 3D deformation integrates simulation with neural representations. Material Point Method (MPM) [SSC13] handles large deformations and topology changes. PhysGaussian [XZQ24] treats 3D Gaussians as simulation particles, extended to language-driven property assignment [QYZW24, ZWZ24b] and text-to-3D synthesis. NeRF methods embed continuum mechanics: PIE-NeRF [FSL24] and PAC-NeRF [LQC23] enable interactive elastodynamics, while Nerfies [PSB21] regularizes deformation with elastic priors. Alternative frameworks explore finite elements, spring-mass systems [ZYWL24], and position-based dynamics [ACRDS23]. Point cloud methods [CZH25b] use elasticity as constraints, while differentiable physics [HZL24, LSY23] leverages neural networks to distill physical laws.
@@ 3.3 Tracking-based Representation
Reliable 2D tracking establishes temporal correspondences foundational for 3D motion recovery. Classical methods include sparse feature matching (e.g., ORB [RRKB11]) and optical flow [LK81, SYLK18, TD20]. Sparse matching supports localization but fails for dense non-rigid motion; optical flow provides dense correspondences but struggles with occlusion and large viewpoint changes. Recent neural rendering methods [WELG21, WCC23, LKLR24] enable dense, long-range 3D tracking by jointly optimizing scene representation and a continuous trajectory field:
@@
Despite its accessibility, 2D tracking faces several critical challenges for 4D tasks: (a)Depth ambiguity: 2D projections lose 3D geometric information, making it difficult to distinguish between different 3D motions that produce similar 2D displacement textit(b) Occlusion handling: Points disappear and reappear in 2D views, breaking temporal correspondence (c) Perspective distortions: 2D tracking doesn’t capture actual 3D motion—camera viewpoint changes confound object motion (d) Requires lifting: An additional step is needed to reconstruct 3D/4D information from 2D observations, introducing errors and ambiguity
@@
3.3.2 Tracking in 3D
The 3D scene flow provides a compact representation of frame-to-frame motion in 4D reconstruction. In the continuous setting, a velocity field $\mathbf{v}(\mathbf{x},t)$ assigns each point a motion vector describing its instantaneous change [NMOG19, ZSX24, WMJL23], where a point’s new position can be approximated as
@@ 3.3.2 Tracking in 3D
However, scene flow methods have notable limitations: they are sensitive to occlusions and visibility constraints, leading to tracking failures or outliers in stereo matching [CHK23]. Highly deformable or low-texture scenes detoriates self-occlusion and reduce accuracy [DMY23a], while frame-wise processing struggles with topology changes and temporal inconsistency. Moreover, limited sensor visibility restricts motion estimation to observed surfaces, leaving discontinuities in the reconstructed motion field [LTT21].
@@ 3.3.3 Integrating physical priors
Integrating physics into 2D tracking helps regularize the inherently ill-posed monocular 4D reconstruction problem through motion constraints and physical priors. Optical flow-based methods [WYG24, GXC24] combine flow and depth cues with physical motion models to enforce coherent spatio-temporal dynamics. Deformation-based approaches apply physics-inspired constraints such as as-rigid-as-possible (ARAP) regularization [LWH24] to preserve local rigidity when lifting 2D priors to 3D. Meanwhile, learning-based frameworks [FZW25, NMOG19] exploit 2D correspondences and monocular depth through reprojection or implicit dynamic modeling, achieving 4D reconstruction without explicit supervision.
Physics integration in 3D tracking depends on the choice of 3D representation. For rigid-body/mesh models, physics-based filters like [Y23] embed differentiable simulators into Extended Kalman Filters (EKF) to track pose, velocity, and friction, while [B21] applies differentiable EKFs to simpler 1D sliding motion. Gaussian/particle representations couple dynamics with rendering: [LKLR24] models scenes as moving 3D Gaussians with local rigidity constraints, [Z24] learns dynamics via Graph Neural Networks on sparse particles, and [X24] enriches Gaussians with Newtonian mechanics (strain, stress, inertia). For extended object tracking[L19] uses Gaussian processes to jointly infer shape and kinematics, while [C24] employs object-centric Gaussian splatting to reconstruct and track dynamic objects in RGB-D sequences.
@@ 3.4 Hybrid Representation
where $\mathcal{R}_{\theta}$ is the rendering function, $\tau(g,t)$ captures geometric variations such as 4D $\sum,\sigma$ in Gaussian Splatting or color density in NeRFs, and $t$ denotes direct time input (e.g., frame index or learnable per-frame latents). The key design choice lies in modeling $\tau(g,t)$: feedforward models [XLD25, MCY25] decode per-pixel 4D Gaussian vectors via DiT heads, while [DWD24, YPZ24] instead use 1D temporal Gaussians. NeRF-based approaches progress from MLPs to 4D neural voxels for faster rendering [PSJ23, GXH23]. In all cases, motion is implicitly embedded in the scene’s geometry and supervised purely from 2D images, making this paradigm computationally efficient and well suited to recent feedforward 4D generation models [MCY25, XLD25].
Per-frame modeling Per-frame 4D reconstruction methods treat each timestep as an independent or weakly coupled 3D reconstruction task [WZH25], producing discrete per-frame 3D representations that are subsequently assembled into a 4D sequence. L4GM [RXM24] generates frame-wise Gaussian representations from LGM [TCC24] conditioned on predictions from previous timesteps, while point-based approaches [WZH25, CCX25a, ZHH24] adopt DusT3R [WLC24] independently for each frame to obtain 4D outputs. However, this frame-by-frame generation strategy typically necessitates interpolation or post-processing to enforce temporal smoothness. For example, [RXM24] employs a video interpolation model [BRL23] to refine renderings, [WZH25, ZHH24] apply 2D point tracking, and [CCX25a] incorporates temporal attention modulation to enhance overall reconstruction fidelity.
@@ 4 Modeling Interaction
Thus far, we have focused on modeling the geometry and motion of 4D entities, in isolation. It is often necessary, however, to model the interactions between multiple 4D entities: between human and human, human and object, or object and object. Such interactions can be learned either from data priors [LHB25] where knowledge of dynamics and causality is implicitly captured from large-scale observations, or from physical simulations [LKJ20] where interactions are explicitly governed by forces, materials, and physics-based rules. We refer readers to surveys dedicated to these two perspectives [ZMR23, SGH25, FHCD25]. In this section, we take a complementary angle, surveying explicit interaction representations that describe the structure of interactions, independent of specific objects or scenes (Figure˜4). We begin by surveying the overall representation strategies for entities involved in interactions. We then delve deeper into three aspects crucial in interaction: pose, contact, action and affordance. Finally, we examine strategies for integrating physical priors into interaction representations.
@@ 4.1 Representing Interaction Entities
where each $M_{i}=\{\mathbf{x}_{i}^{1},\mathbf{x}_{i}^{2},\ldots,\mathbf{x}_{i}^{T}\}$ is a sequence of states $x_{i}^{t}$ describing the motion of entity $E_{i}$ over time steps $1\ldots T$. The representation choice of each entity depends on both its individual properties, as well as the property of the entire interaction. Humans are most commonly represented with skeletal poses or template meshes (Section 2.2.1), as these excel at modeling long motion sequences which are central to most interaction events. Given the complexity of interaction data, simpler representations are sometimes favored to make the problem tractable. To address the weaknesses of template meshes, a growing trend is to couple them with neural field based representations to obtain increased geometric fidelity [PSX24, QLC25] and more flexible, differentiable optimization of poses and contact objectives [OGK25, WHZ25a, LYLW24, CPH24]. For other objects, meshes (Section 2.1.1) are the popular choice due to having explicit surface representations, which enables more efficient modeling of contact and physics. Despite these advances, the assumption of object rigidity remains a major simplification. Extending to articulated objects brings more flexibilty, but also introduces the need for a unified representation that captures part connectivity, motion patterns, and pose variations beyond fixed templates.
@@ 4.3 Contact
Interaction entities contact with each other. Accurately representing contact can either be an end goal, or a means for more accurate interaction modeling. Explicit representations of contact, such as contact maps [BHKH19, BTT20, TGBT20] and neural distance fields, have been shown to improve motion generation. [MHS22] predicts future contact maps enhances both local and global motion consistency. GanHand [CPA20] refines contact modeling at finger and force levels to improve pose estimation. Neural distance fields (NDF) emerges as an auxiliary contact representation. NDF [KYZ20, TAL22, WHMM22, UFPC22, KRG24] expand the single-frame local contact information into a spatially-dense field to guide motion generation. Grasping Fields [KYZ20] encodes the distances to valid grasps, and PoseNDF [TAL22] learns encode full-body pose into the field. NIFTY [KRG24] encode the body pose gradient towards interacting frame into the fields. All these works demonstrate that explicit contact modeling improves generation. Most existing research, however, relies on SMPL or MANO, which are rigid representation and naturally induces penetration. Another future step might be to extend contact modeling beyond the visual domain to incorporate mechanical realism and force reasoning.
@@ 5 Datasets and Benchmarks
Figure 5: Overview of key datasets in 4D representation research: (a) Objaverse-Dy [YXV25] provides a large-scale collection of synthetic animated 3D assets, serving as the primary source for current 4D generation works after filtering; (b) ActorsHQ [IRG23] is a high-fidelity, multi-view human performance capture dataset with per-frame 3D mesh reconstructions for dynamic human motion (c) Stereo4D [JTL24] represent emerging large-scale datasets capturing native 4D data from real-world scenarios.(d) PartNet-Mobility [XQM20] offers an articulated object database that facilitates part-based representation and articulated motion learning; (e) Ego-Exo4D [GWT24] is a multi-camera synchronized video corpus capturing first-person (egocentric) and third-person (exocentric) views of skilled human activities. (f) Parahome [KKNJ25] is a dataset capturing 3D human–object interaction in a natural home environment and (g) DeformingThings4D [LTT21] is a large-scale synthetic dataset of non-rigidly deforming 3D objects (humanoids and animals) with dense 4D annotations; Actors-HQ, Stereo4D and Ego-Exo4D represent emerging large-scale datasets capturing native 4D data from real-world scenarios.
Unlike other computer vision tasks with abundant labeled data, 4D representation learning faces a data scarcity problem. The ideal dataset would contain real-world dynamic scenes with complete ground-truth geometry, appearance, motion and interaction annotations. However, such datasets remain limited due to technical and economic challenges in capturing high-quality spatio-temporal data. To resolve this issue, current 4D field leverages a mixture of dataset ranging from 2D to 4D dimensions to guide 4D representation learning ( Figure˜5). This section surveys the current dataset in use from the perspective of learning geometry, motion and interactions.
@@ 5.1 Geometry Datasets
Geometric learning in 4D works primarily leverages two data paradigms: static 3D object datasets for feed-forward reconstruction and large-scale 2D image collections for score distillation-based generation.
Static 3D Object Datasets. Modern 4D generation systems [CZTW25, SLY25] employ feed-forward 3D generation models [XLX25, LZL25, ZLL25] to first synthesize static geometry, then animate it through video guidance or rigging techniques. These feed-forward models are enabled by large-scale 3D object collections. Primary datasets include ShapeNet [CFG15] (51K objects) and ModelNet [WSK15] (12K objects), while Objaverse-XL [DLW23] dramatically scales to 10+ million models, becoming one of the most widely adopted datasets in current research. However, Objaverse exhibits significant quality variance, allowing researchers to query high-quality filtered subsets [XLX25, ZWZ24a]. Other smaller high-quality 3D collections including ABO [CGD22], 3D-Future [FJG21], OmniObject3D [WZF23], Toy4K [STR21], HSSD [KMJ24], and GSO [DFK22] are commonly used as training set additives or benchmark dataset. These datasets provide native supervision for high-quality 3D reconstruction.
@@ 5.2 Motion Datasets
Large-Scale Video Collections. Video diffusion model is widely used modules guide the motion synthesis in current 4D works [LZZ24, LZRV24b, VNZ25, JYC24], while images can only capture static geometry features, videos can jointly capture the geometry and motion. WebVid-10M [BNVZ21] and HD-VILA-100M [XHZ22] stands for two of the large-scale video dataset. These collections enable models like Align Your Latents [BRL23] and Make-a-Video [SPH22], which ensbled methods such as  [SSP23, LKT24] that uses video diffusion model into 4D generation. Adopting a similar strategy to improve image diffusion models, the video diffusion models are also finetuned [XYV24, YXV25, WYWB25, LYX24a] on filtered dynamic objects from Objaverse to improve on multiview motion-consistency in video generations. The primary advantage of applying video diffusion models lies in its generalizability learned from immersive data, supporting motion synthesis across varied applications. However, 2D video supervision prioritizes frame-level realism over 3D motion constraints, often producing sequences that appear locally realistic but violate physical continuity when viewed as coherent 3D trajectories.
@@
Specialized Multi-View Captures. Domain-specific datasets provide high-fidelity temporal ground truth through controlled multi-view recording systems. For human performance, representative datasets include ZJU-MoCap [PZX21], DynaCap [HLX21], ActorsHQ [IRG23], and Human3.6M [IPOS13]. For general dynamic scenes, D-NeRF [PCPMMN21], Nerfies [PSB21], HyperNerf [PSH21], and Plenoptic Video [LSZ22] capture real-world deformations with varying complexity. These datasets preserve authentic physical dynamics and provide strong multi-view supervision for consistent 3D motion reconstruction. However, limited scale—typically dozens to hundreds of sequences compared to millions in large-scale video collections—and substantial capture infrastructure requirements constrain their broader applicability.
@@
Animated Motion Libraries. Synthetic motion captures offer perfect ground truth with structural priors for controllable animation. AMASS [MGT19] consolidates over 40 hours of motion capture from 15 datasets into unified SMPL [LMR15] parametric representations for human motion synthesis. Mixamo [Ado15] provides thousands of professionally rigged character animations, while Truebones Zoo [Tru] extends coverage to 75+ non-humanoid creatures. For animated objects, curated Objaverse subsets [LYX24a, XYV24, LZRV24b, LZZ24] filter hundreds of thousands of animated assets based on motion quality and temporal coherence, with Articulation-XL [XLX25] providing 48K+ models featuring automated skeleton quality assessment via Vision-Language Models. These synthetic collections enable large-scale training with explicit motion annotations, though they exhibit domain gaps and simplified physics compared to real-world capture, particularly for complex interactions and environmental effects.
@@
Deformable object collections provide ground truth for non-rigid motion learning. DeformingThings4D [LTT21] contains 1,972 synthetic animation sequences spanning 31 categories (humanoids and animals) with dense 4D annotations including signed distance fields, volumetric motion fields, and scene flow at multiple hierarchical resolutions. The most recent effort, Stereo4D [JTL24] offers 100K+ real-world clips mined from Internet VR180 videos, providing pseudo-metric 3D point clouds with long-range world-space trajectories derived through careful fusion of stereo depth estimation, 2D tracking, and structure-from-motion. These datasets address the scarcity of real-world 4D supervision—DeformingThings4D through high-quality synthetic rendering and Stereo4D through scalable data.

### v2

@@ 3.2.1 Rigid Motion
Rigid bodies include manufactured items (boxes, chairs, kitchenware) and vehicles, which are often approximated as rigid despite minor internal deformations [PYL22, YAK20, XH22].
@@ 3.2.2 Non-Rigid Motion
where $\Delta_{\theta}$ predicts displacements conditioned on temporal code $\tau(t)$. These fields are trained from 2D observations via differentiable rendering, eliminating 3D supervision. However, sparse views complicate learning, requiring regularization for robustness.
@@ 3.2.3 Integrating physical priors
Physics-based 3D deformation integrates simulation with neural representations. Material Point Method (MPM) [SSC13] handles large deformations and topology changes. PhysGaussian [XZQ24] treats 3D Gaussians as simulation particles, which has been extended to language-driven property assignment [QYZW24] and text-to-3D synthesis. NeRF methods embed continuum mechanics: PIE-NeRF [FSL24] and PAC-NeRF [LQC23] enable interactive elastodynamics, while Nerfies [PSB21] regularizes deformation with elastic priors. Alternative frameworks explore finite elements, spring-mass systems [ZYWL24], and position-based dynamics [ACRDS23].  [CZH25] uses elasticity as a constraint in point cloud modelling, while differentiable physics [HZL24, LSY23] leverages neural networks to distill physical laws.
@@ 3.3 Tracking-based Representation
Reliable 2D tracking establishes temporal correspondences foundational for 3D motion recovery. Classical methods include sparse feature matching (e.g., ORB [RRKB11]) and optical flow [SYLK18, TD20]. Sparse matching supports localization but fails for dense non-rigid motion; optical flow provides dense correspondences but struggles with occlusion and large viewpoint changes. Recent neural rendering methods [WELG21, WCC23, LKLR24] enable dense, long-range 3D tracking by jointly optimizing scene representation and a continuous trajectory field:
@@
Despite its accessibility, 2D tracking faces several critical challenges for 4D tasks: (a) Depth ambiguity: 2D projections lose 3D geometric information, making it difficult to distinguish between different 3D motions that produce similar 2D displacement. (b) Occlusion handling: Points disappear and reappear in 2D views, breaking temporal correspondence. (c) Perspective distortions: 2D tracking doesn’t capture actual 3D motion—camera viewpoint changes confound object motion. (d) Requires lifting: An additional step is needed to reconstruct 3D/4D information from 2D observations, introducing errors and ambiguity.
@@
3.3.2 3D Tracking
The 3D scene flow provides a compact representation of frame-to-frame motion in 4D reconstruction. In the continuous setting, a velocity field $\mathbf{v}(\mathbf{x},t)$ assigns each point a motion vector describing its instantaneous change [NMOG19, WMJL23], where a point’s new position can be approximated as
@@
However, scene flow methods have notable limitations: they are sensitive to occlusions and visibility constraints, leading to tracking failures [LZYX22]. Highly deformable or low-texture scenes exacerbate self-occlusion and reduce accuracy [DMY24], while frame-wise processing struggles with topology changes and temporal inconsistency. Moreover, limited sensor visibility restricts motion estimation to observed surfaces, leaving discontinuities in the reconstructed motion field [LTT21].
@@ 3.3.3 Integrating physical priors
Integrating physics into 2D tracking regularizes the ill-posed monocular reconstruction problem through motion constraints and physical priors. Optical-flow-based methods [WYG24, GXC24] combine flow and depth cues with physical motion models to enforce coherent spatio-temporal dynamics. Deformation-based approaches apply physics-inspired constraints including as-rigid-as-possible (ARAP) regularization [LWH25] to preserve local rigidity. Meanwhile, learning-based frameworks [FZW25, NMOG19] exploit 2D correspondences and monocular depth through reprojection or implicit dynamic modeling, achieving 4D reconstruction without explicit supervision.
Physics integration in 3D tracking depends on the choice of 3D representation. For rigid-body/mesh models, physics-based filters like [KSS24] embed differentiable simulators into Extended Kalman Filters (EKF) to track pose, velocity, and friction, while [PPN21] applies differentiable EKFs to simpler 1D sliding motion. Gaussian/particle representations couple dynamics with rendering: [LKLR24] models scenes as moving 3D Gaussians with local rigidity constraints, [ZZL24] learns dynamics via Graph Neural Networks on sparse particles, and [XZQ24] enriches Gaussians with Newtonian mechanics (strain, stress, inertia). For extended object tracking[KO21] uses Gaussian processes to jointly infer shape and kinematics, while [BRD24] employs object-centric Gaussian splatting to reconstruct and track dynamic objects from RGBD.
@@ 3.4 Hybrid Representation
where $\mathcal{R}_{\theta}$ is the rendering function, $\tau(g,t)$ captures geometric variations such as 4D $\sum,\sigma$ in Gaussian Splatting or color density in NeRFs, and $t$ denotes direct time input (e.g., frame index or learnable per-frame latents). The key design choice lies in modeling $\tau(g,t)$: feed-forward models [XLD25, MCY25] decode per-pixel 4D Gaussian vectors via DiT heads, while [DWD24, YPZ24] instead use 1D temporal Gaussians. NeRF-based approaches progress from MLPs to 4D neural voxels for faster rendering [PSJ23, GXH23]. In all cases, motion is implicitly embedded in the scene’s geometry and supervised purely from 2D images, making this paradigm computationally efficient and well-suited to recent feed-forward 4D generation models [MCY25, XLD25].
Per-frame modeling Per-frame 4D reconstruction methods treat each timestep as an independent or weakly coupled 3D reconstruction task [WZH25], producing discrete per-frame 3D representations that are subsequently assembled into a 4D sequence. L4GM [RXM24] generates frame-wise Gaussian representations from LGM [TCC24] conditioned on predictions from previous timesteps, while point-based approaches [WZH25, CCX25a, ZHH24] adopt DusT3R [WLC24] independently for each frame to obtain 4D outputs. However, this frame-by-frame generation strategy typically necessitates interpolation or post-processing to enforce temporal smoothness. For example, [RXM24] employs a video interpolation model [BRL23] to refine renderings, and [CCX25a] incorporates temporal attention modulation to enhance overall reconstruction fidelity.
@@ 4 Modeling Interaction
Thus far, we have focused on modeling the geometry and motion of 4D entities, in isolation. It is often necessary, however, to model the interactions between multiple 4D entities: between human and human, human and object, or object and object. Such interactions can be learned either from data priors [LHB25] where knowledge of dynamics and causality is implicitly captured from large-scale observations, or from physical simulations [LKJ20] where interactions are explicitly governed by forces, materials, and physics-based rules. We refer readers to surveys dedicated to these two perspectives [ZMR23, SGH26, FHCD25]. In this section, we take a complementary angle, surveying explicit interaction representations that describe the structure of interactions, independent of specific objects or scenes (Figure˜4). We begin by surveying the overall representation strategies for entities involved in interactions. We then delve deeper into three aspects crucial in interaction: pose, contact, action and affordance. Finally, we examine strategies for integrating physical priors into interaction representations.
@@ 4.1 Representing Interaction Entities
where each $M_{i}=\{\mathbf{x}_{i}^{1},\mathbf{x}_{i}^{2},\ldots,\mathbf{x}_{i}^{T}\}$ is a sequence of states $x_{i}^{t}$ describing the motion of entity $E_{i}$ over time steps $1\ldots T$. The representation choice of each entity depends on both its individual properties, as well as the property of the entire interaction. Humans are most commonly represented with skeletal poses or template meshes (Section 2.2.1), as these excel at modeling long motion sequences which are central to most interaction events. Given the complexity of interaction data, simpler representations are sometimes favored to make the problem tractable. To address the weaknesses of template meshes, a growing trend is to couple them with neural field based representations to obtain increased geometric fidelity [PSX24, QLC25] and more flexible, differentiable optimization of poses and contact objectives [OGK25, WHZ25, LYLW24, CPH24]. For other objects, meshes (Section 2.1.1) are the popular choice due to having explicit surface representations, which enables more efficient modeling of contact and physics. Despite these advances, the assumption of object rigidity remains a major simplification. Extending to articulated objects brings more flexibility, but also introduces the need for a unified representation that captures part connectivity, motion patterns, and pose variations beyond fixed templates.
@@ 4.3 Contact
Interacting entities make contact with each other. Accurately representing contact can either be an end goal, or a means for more accurate interaction modeling. Explicit representations of contact, such as contact maps [BHKH19, BTT20, TGBT20] and neural distance fields, have been shown to improve motion generation. [MHS22] shows that predicting future contact maps enhances both local and global motion consistency. GanHand [CPA20] refines contact modeling at finger and force levels to improve pose estimation. Neural distance fields (NDF) emerges as an auxiliary contact representation. NDF [KYZ20, TAL22, WHMM22, UFPC22, KRG24] expands the single-frame local contact information into a spatially-dense field to guide motion generation. Grasping Fields [KYZ20] encodes the distances to valid grasps, and PoseNDF [TAL22] learns to encode full-body pose into the field. NIFTY [KRG24] encodes the body pose gradient towards interacting frame into the fields. All these works demonstrate that explicit contact modeling improves generation. Most existing research, however, relies on SMPL or MANO, which are rigid representations and naturally induce penetration. Another future step might be to extend contact modeling beyond the visual domain to incorporate mechanical realism and force reasoning.
@@ 5 Datasets and Benchmarks
Figure 5: Overview of key datasets in 4D representation research: (a) Objaverse-Dy [YXV25] provides a large-scale collection of synthetic animated 3D assets, serving as the primary source for current 4D generation works after filtering; (b) Actors-HQ [IRG23] is a high-fidelity, multi-view human performance capture dataset with per-frame 3D mesh reconstructions for dynamic human motion; (c) Stereo4D [JTL24] represents emerging large-scale datasets capturing native 4D data from real-world scenarios; (d) PartNet-Mobility [XQM20] offers an articulated object database that facilitates part-based representation and articulated motion learning; (e) Ego-Exo4D [GWT24] is a multi-camera synchronized video corpus capturing first-person (egocentric) and third-person (exocentric) views of skilled human activities; (f) Parahome [KKNJ25] is a dataset capturing 3D human–object interaction in a natural home environment, and (g) DeformingThings4D [LTT21] is a large-scale synthetic dataset of non-rigidly deforming 3D objects (humanoids and animals) with dense 4D annotations.
Unlike other computer vision tasks with abundant labeled data, 4D representation learning faces a data scarcity problem. Ideal datasets would contain real-world dynamic scenes with complete ground-truth geometry, appearance, motion and interaction annotations. However, such datasets remain limited due to technical and economic challenges in capturing high-quality spatio-temporal data. To address this, the current 4D field leverages a mixture of datasets ranging from 2D to 4D dimensions to guide 4D representation learning  (Figure˜5). This section surveys the current dataset in use from the perspective of learning geometry, motion and interactions.
@@ 5.1 Geometry Datasets
Geometric learning in 4D works primarily leverages two data paradigms: static 3D object datasets for feed-forward reconstruction and large-scale 2D image collections for score-distillation-based generation.
Static 3D Object Datasets. Modern 4D generation systems [CZTW25, SLY25] employ feed-forward 3D generation models [XLX25, LZL25, ZLL25] to first synthesize static geometry, then animate it through video guidance or rigging techniques. These feed-forward models are enabled by large-scale 3D object collections. Primary datasets in early time include ShapeNet [CFG15] (51K objects) and ModelNet [WSK15] (12K objects), while the more recent Objaverse-XL [DLW23] dramatically scales to 10+ million models, becoming one of the most widely adopted datasets in current research. However, Objaverse exhibits significant quality variance, leading researchers to create high-quality filtered subsets [XLX25, ZWZ24]. Other smaller high-quality 3D collections, including ABO [CGD22], 3D-Future [FJG21], OmniObject3D [WZF23], Toy4K [STR21], HSSD [KMJ24], and GSO [DFK22] are commonly used as training set additives or benchmark datasets. These datasets provide native supervision for high-quality 3D reconstruction.
@@ 5.2 Motion Datasets
Large-Scale Video Collections. Video diffusion models serve as key components for guiding motion synthesis in 4D works [LZZ24, LZRV24b, VNZ25, JYC24]. Unlike images that capture only static geometry, videos encode both geometric structure and temporal dynamics. WebVid-10M [BNVZ21] and HD-VILA-100M [XHZ22] represent two prominent large-scale video datasets that enable training models like Align Your Latents [BRL23] and Make-a-Video [SPH23]. These foundational video diffusion models have been successfully adapted for 4D generation [SSP23, LKT24]. Following strategies used to enhance image diffusion models, recent work finetunes video diffusion models [XYV24, YXV25, WYWB25, LYX24a] on filtered dynamic objects from Objaverse, significantly improving multiview motion consistency. The primary advantage of video diffusion models lies in their generalizability from large-scale training data, enabling motion synthesis across diverse applications. However, 2D video supervision prioritizes frame-level visual realism over 3D geometric constraints, often producing sequences that appear locally plausible but violate physical continuity when reconstructed as coherent 3D trajectories.
@@
Specialized Multi-View Captures. Domain-specific datasets provide high-fidelity temporal ground truth through controlled multi-view recording systems. For human motion, representative datasets include ZJU-MoCap [PZX21], DynaCap [HLX21], ActorsHQ [IRG23], and Human3.6M [IPOS13]. For general dynamic scenes, D-NeRF [PCPMMN21], Nerfies [PSB21], HyperNerf [PSH21], and Plenoptic Video [LSZ22] capture real-world deformations with varying complexity. These datasets preserve authentic physical dynamics and provide strong multi-view supervision for consistent 3D motion reconstruction. However, limited scale—typically dozens to hundreds of sequences compared to millions in large-scale video collections—and substantial capture infrastructure requirements constrain their broader applicability.
@@
Animated Motion Libraries. Synthetic motion captures offer perfect ground truth with structural priors for controllable animation. AMASS [MGT19] consolidates over 40 hours of motion capture from 15 datasets into unified SMPL [LMR15] parametric representations for human motion synthesis. Mixamo [Ado15] provides thousands of professionally rigged character animations, while Truebones Zoo [Tru] extends coverage to 75+ non-humanoid creatures. For animated objects, curated Objaverse subsets [LYX24a, XYV24, LZRV24b, LZZ24] provide hundreds of thousands of animated assets filtered by motion quality and temporal coherence. Articulation-XL [XLX25] provides 48K+ models featuring automated skeleton quality assessment via Vision-Language Models. These synthetic collections enable large-scale training with explicit motion annotations, though they exhibit domain gaps and simplified physics compared to real-world capture, particularly for complex interactions and environmental effects.
@@
Deformable Object Collections provide ground truth for non-rigid motion learning. DeformingThings4D [LTT21] contains 1,972 synthetic animation sequences spanning 31 categories (humanoids and animals) with dense 4D annotations including signed distance fields, volumetric motion fields, and scene flow at multiple hierarchical resolutions. The most recent effort, Stereo4D [JTL24] offers 100K+ real-world clips mined from Internet VR180 videos, providing pseudo-metric 3D point clouds with long-range world-space trajectories derived through careful fusion of stereo depth estimation, 2D tracking, and structure-from-motion. These datasets address the scarcity of real-world 4D supervision—DeformingThings4D through high-quality synthetic rendering and Stereo4D through scalable data collection.